OpenStudy (babyslapmafro):

Please help me find the anti-derivative of x/(x^2+1)

OpenStudy (babyslapmafro):

I'm using the shell method to find the volume of the solid generated when the region enclosed by \[y=\frac{ 1 }{ x^2+1 },y=0 \] \[x=0,x=1\] is revolved about the y-axis

OpenStudy (babyslapmafro):

this is where i am in the problem, just need some help integrating \[2\pi \int\limits_{0}^{1}\frac{ x }{ x^2+1 }dx\]

terenzreignz (terenzreignz):

Let's write it this way, maybe it'll be easier :) \[\pi \int\limits\limits_{0}^{1}\frac{ 2x }{ x^2+1 }dx\] better?

OpenStudy (babyslapmafro):

hmm...still stuck

terenzreignz (terenzreignz):

lol, if you take that denominator, its derivative is to be found somewhere in the integrand...

OpenStudy (babyslapmafro):

ok, the derivative of the denominator is 2x

terenzreignz (terenzreignz):

Which is the numerator... can you see it now? :)

OpenStudy (babyslapmafro):

ln(x^2+1)

terenzreignz (terenzreignz):

That's good, now evaluate it from x = 0 to x = 1 And then multiply it by pi, and you'll have your result :)

OpenStudy (babyslapmafro):

So multiplying the integrand by two before integrating doesn't effect the final result negatively?

terenzreignz (terenzreignz):

I didn't multiply the integrand by 2, at least, I didn't do JUST that, remember, the number outside the integrand was 2pi? I just brought the 2 inside the integrand (you can do that with constants) so you could see that the derivative of the denominator can be found in the integrand.

OpenStudy (babyslapmafro):

ok thanks